# Predicting Grain Boundary Segregation in Magnesium Alloys: An Atomistically Informed Machine Learning Approach

### v1

@@ IV.1 Signature energetic and structural descriptors
The SOAP-featured ML model demonstrated superior performance in predicting $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$ at 600 K at GB sites in the Mg polycrystal compared to the ESD-featured model. The SOAP descriptors define the atomic density constructed from Gaussians centered on each site, effectively representing the local atomic environment [58, 59]. Although SOAP has been widely used for generating structural features for training ML models to predict not only $\Delta E_{\text{seg}}$ but also other GB properties [66, 67], its limitations include the statistical significance of utilizing a large number of features (over a thousand variables) in the training procedure, which increases the risk of learning spurious correlations, and the lack of physical insight obtained from the approach. In contrast, the ESD feature set consists of important properties of the local atomic environment, each with its physical significance, which can help achieve a better understanding of the dominant properties influencing segregation behavior and develop physics-based ML models.
Among all ESD features, $\sigma_{h}$ stands out as the most important feature in predicting $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$. For solute elements with larger atomic radii than Mg, a negative correlation with target segregation energies was observed, while for solute elements with smaller atomic radii, a positive correlation was identified (see Figure S4-9). Such solute size-dependent segregation behavior has been widely reported in previous experimental and atomic-scale modeling works [11, 32, 13, 14], namely, the preference for large solutes to segregate at sites with extensive strain and small solutes to segregate at sites with compressive strain. Huber et al. [32] developed a linear elastic model for predicting $\Delta E_{\text{seg}}$ based on this concept, where the multiplication of variate in atomic volume and bulk modulus estimates the isotropic stress state at the GB site.
@@
Although the prediction of overall free energy of solute segregation is reliable, the performance of the ML approach in predicting the vibrational entropy contribution $\Delta F_{\text{seg}}^{\text{vib}}$ is unsatisfactory using both ESD and SOAP feature sets. $\Delta F_{\text{seg}}^{\text{vib}}$ is directly related to segregation entropy ($\Delta S_{\text{seg}}$), as demonstrated in Equation 3 and 4. Rittner and Seidman [72] reported a strong linear correlation between $\Delta E_{\text{seg}}$ and $\Delta S_{\text{seg}}$ for Pd solute segregation in symmetric tilt Ni GBs using atomistic simulations. The experimental data on Si, P, and C solute segregation in tilt GBs in Fe-Si alloy systems from Lejček and Hofmann [73] also demonstrated a linear segregation energy-entropy relationship. Tuchinda and Schuh [57] reported the linear segregation energy-entropy relationship in the Ni-based polycrystalline alloys using atomistic simulations and attributed it to the physical principle that deeper energy wells are narrower, thus restricting vibration. However, significant scattering was detected and fitted to a bivariate spectral isotherm [57]. Similar bivariate skew-normal distributions were also identified in the Mg alloying systems in this study, as illustrated in Figure S34. A linear function was fitted to the distribution, with the slope having a similar magnitude ($10^{-4}$ 1/K) as that obtained for the Ni-based alloy systems [57]. Such significant scattering between $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}^{\text{vib}}$ leads to the unsatisfactory prediction of $\Delta F_{\text{seg}}^{\text{vib}}$ using the $\Delta E_{\text{seg}}$ feature. The prediction becomes even worse without including $\Delta E_{\text{seg}}$ in the feature sets, as $\Delta E_{\text{seg}}$ exhibits the highest correlation with $\Delta F_{\text{seg}}^{\text{vib}}$ among all features, as illustrated in the correlation matrices (Figure S4-9) and feature importance (Figure S15). The SOAP feature set, which is believed to comprehensively capture the structural features of the local atomic environment, fails to predict $\Delta F_{\text{seg}}^{\text{vib}}$, implying that structural features before solute segregation alone are insufficient for predicting vibrational entropy contribution of solute segregation. This can be interpreted as the change in bond stiffness, crucial in determining the variation of vibrational frequency when introducing a solute at a GB site, which cannot be precisely described using the structural features of the pre-segregation local atomic environment. Other energetic features describing the system before solute segregation, such as potential energy and $F^{\text{vib}}$, although still among the highest important features, exhibit much lower importance compared to $\Delta E_{\text{seg}}$ (Figure S15). Overall, these outcomes highlight that the features after segregation are crucial in predicting the vibrational entropy contribution, with $\Delta E_{\text{seg}}$ being particularly important as it describes the energy difference before and after introducing the solute element. For better prediction of $\Delta F_{\text{seg}}^{\text{vib}}$, incorporating more structural and energetic features after solute segregation will be beneficial.
@@
Figure 7: (a) Summary of experimentally measured concentrations of various solutes at GB versus bulk in Mg alloys [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. The concentrations of (b) Nd, (c) Ca, (d) Y, (e) Al, and (f) Zn solutes at GBs within the simulated Mg polycrystal as a function of their bulk solute concentrations, incorporating the vibrational entropy contribution $\Delta E_{\text{seg}}+\Delta F_{\text{seg}}^{\text{vib}}$ (represented by solid lines) and considering only the segregation energy $\Delta E_{\text{seg}}$ (represented by dashed lines) across various temperature regimes following the Langmuir–McLean model, as denoted in Equations 12 and 13. Experimental data presented in (a) are illustrated as black dots on the corresponding graphs, labeled with the first authors’ names and the thermomechanical processing temperatures of the Mg alloys.
@@
Experimental data of $X_{\text{GB}}$ versus $X_{\text{bulk}}$ are summarized in Figure 7(a). Notably, the Nd solute exhibits significantly higher $X_{\text{GB}}$ than other solute elements under similar $X_{\text{bulk}}$ and temperature conditions. This experimental observation is in good agreement with the simulation results presented in Figure 3, which demonstrate an elevated $X_{\text{GB}}$ for Nd at both 0.1 at.and 1 at.$X_{\text{bulk}}$ across the temperature range up to 800 K. This consistency between experimental and simulated data underscores Nd’s exceptional energetical favorability for segregation at GBs, distinguishing it as the most prone among the alloying elements studied to segregate. However, it is important to note that the simulation predicts the $X_{\text{GB}}$ of Nd solute with an overestimation by a factor of 4 to 6 in the temperature regime from 500 to 600 K (see Figure 7(b)). An overestimation in $X_{\text{GB}}$ for the Ca and Al solutes was also identified, with a factor of 2 to 3 across the relevant temperature and $X_{\text{bulk}}$ regimes (see Figure 7(c,e)). A relatively good agreement was observed between experimental measurements and simulated predictions of $X_{\text{GB}}$ for Y and Zn solutes. The experimentally measured $X_{\text{GB}}$ values fall within the range predicted by atomistic simulations. No experimental data regarding the chemical distribution of Li in GB segregation were available to the best of our knowledge, primarily due to the research focus on forming Mg-Li intermetallic phases to enhance the mechanical performances of Mg alloys rather than on engineering the GB properties below the solid solubility limit [74, 75].
@@
The overestimation of $X_{\text{GB}}$ could be attributed to factors related to both theoretical methodologies and experimental measurements. Firstly, the Langmuir–McLean model, employed in our current study, assumes an infinitesimally diluted solid solution, where solute-solute interactions are not accounted for. Experimental Mg alloy systems are often more complex than simple binary compositions, making it necessary to account for synergistic effects from solute-solute interactions. This simplification could contribute to discrepancies between experimental observations and theoretical predictions, e.g., as demonstrated in our recent APT study on the Mg-Zn-Ca alloy systems, the presence of Zn significantly reduced the concentration of Ca at GBs from 2.39 at.in a Mg-Ca binary system where $X_{\text{bulk}}$ of Ca at 0.163 at.to 1.95 at.in a Mg-Zn-Ca ternary system with even higher bulk Ca concentration of 0.421 at.(see Mouhib, 673 K in Figure 7(c)) [31]. Given that this binary alloy system for Ca segregation is the only example available in the literature, it is noteworthy that the experimentally measured $X_{\text{GB}}$ for this system demonstrates a closer alignment with theoretical predictions compared to results from ternary or quaternary alloy systems. Moreover, the available experimental data on Y segregation at GBs were all obtained from Mg-Y binary alloy systems (Figure 7(d)), demonstrating a good correlation with theoretical predictions. To investigate solute segregation behavior at higher concentrations, hybrid Monte-Carlo/molecular dynamics (MC/MD) approaches can be employed to provide thermodynamic insights into solute clustering, precipitation, and the formation of intermetallic phases at GBs [76, 77].
@@
Additionally, the discrepancy may stem from the limitations in experimental measurements, which were performed on selected individual GBs, contrasting with atomistic simulations that encompassed a polycrystal with over 38,000 GB sites in a GB network with various macro and microscopic characteristics. As revealed in our recent APT study, significant variations were observed in the segregation behavior of Nd among GBs in Mg alloys, which possess distinct macroscopic characteristics [30]. Notably, the concentration of Nd at GBs ($X_{\text{GB}}$) was found to be more than twice as high in a GB with a misorientation angle of 61.1 compared to one with an angle of 35.8. This scattering in GB segregation can primarily be attributed to variations in the local atomic environments, indicating the crucial role of microscopic characteristics in influencing solute distribution at the atomic level. Furthermore, in experimental conditions, there is always competition with other defects, such as dislocations or stacking faults, which cannot be switched off. If a large number of solutes get trapped in these defects, this can also lead to less solute being available at the GBs. This underscores the need for defect phase diagrams [78] to better understand solute distribution mechanisms under different chemical potential conditions.
@@
Furthermore, artifacts and approximations in both experimental and simulation methodologies can significantly impact the outcome solute segregation behavior. For instance, during the post-processing of APT and EDS data, assumptions about the GB thickness can lead to notable discrepancies in the measurements of $X_{\text{GB}}$. Similarly, the definition of GB sites according to structural analysis algorithms for segregation energy calculations influences the segregation energy spectrum and, consequently, the estimated solute concentration at GBs. In atomistic simulations, the choice of interatomic potentials and methodological assumptions for calculating $\Delta F_{\text{seg}}^{\text{vib}}$, such as those made by the HA method, can influence results. Specifically, the HA method’s inability to capture anharmonic behaviors—–like structural transitions—–and its omission of other significant entropy contributions, especially at elevated temperatures, are notable limitations of this model. The limitations of semi-empirical potentials in accurately capturing the variation in solute segregation behaviors across the diverse local atomic environments of GBs and a broad range of alloy systems have been reported [79]. Hybrid multiscale modeling approaches, such as quantum mechanical/molecular mechanical (QM/MM), could offer QM-level accurate results while circumventing the size constraints typically associated with ab-initio calculations.
@@ V Conclusions
In this study, the GB segregation behavior of Nd, Ca, Y, Li, Al, and Zn solute elements in polycrystalline Mg at 0 K and finite temperatures was systematically investigated using atomistic simulations. Machine learning (ML) models were employed to predict segregation energy and free energy, incorporating energetic and structural descriptors. The GB solute concentrations in these Mg alloy systems were calculated following the Langmuir–McLean isotherm and compared with experimentally measured local chemical distributions at various temperature and bulk concentration regimes. Key insights from the study are summarized as follows:

### v2

@@ IV.1 Signature energetic and structural descriptors
The SOAP-featured ML model demonstrated superior performance in predicting $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$ at 600 K at GB sites in the Mg polycrystal compared to the ESD-featured model. The SOAP descriptors define the Gaussian smeared atomic density constructed on each site, effectively representing the local atomic environment [61, 62]. While SOAP has been widely utilized for generating structural features to train ML models for predicting not only $\Delta E_{\text{seg}}$ but also other GB properties [71, 72], it has certain limitations. These include the statistical significance of using a large number of features (over a thousand variables) in the training procedure, which raises the risk of learning spurious correlations, and the absence of physical insights obtained from this approach. In contrast, the ESD feature set comprises important properties of the local atomic environment, each with its own physical significance. This can aid in gaining a deeper understanding of the dominant factors influencing segregation behavior and in developing physics-based ML models.
Among all ESD features, $\sigma_{h}$ stands out as the most important feature in predicting $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}$. For solute elements with larger atomic radii than Mg, a negative correlation with target segregation energies was observed, while for solute elements with smaller atomic radii, a positive correlation was identified (see Figures S4-9). Such solute size-dependent segregation behavior has been widely reported in previous experimental and atomic-scale modeling works [16, 37, 18, 19], namely, the preference for large solutes to segregate at sites with extensive strain and small solutes to segregate at sites with compressive strain. Huber et al. [37] developed a linear elastic model for predicting $\Delta E_{\text{seg}}$ based on this concept, where the multiplication of variate in atomic volume and bulk modulus estimates the isotropic stress state at the GB site.
@@
Although the prediction of overall free energy of solute segregation is reliable, the performance of the ML approach in predicting the vibrational entropy contribution $\Delta F_{\text{seg}}^{\text{vib}}$ is unsatisfactory using both ESD and SOAP feature sets. $\Delta F_{\text{seg}}^{\text{vib}}$ is directly related to segregation entropy ($\Delta S_{\text{seg}}$), as demonstrated in Equation (3) and (4). Rittner and Seidman [77] reported a strong linear correlation between $\Delta E_{\text{seg}}$ and $\Delta S_{\text{seg}}$ for Pd solute segregation in symmetric tilt Ni GBs using atomistic simulations. The experimental data on Si, P, and C solute segregation in tilt GBs in Fe-Si alloy systems from Lejček and Hofmann [78] also demonstrated a linear segregation energy-entropy relationship. Tuchinda and Schuh [60] reported the linear segregation energy-entropy relationship in the Ni-based polycrystalline alloys using atomistic simulations and attributed it to the physical principle that deeper energy wells are narrower, thus restricting vibration. However, significant scattering was detected and fitted to a bivariate spectral isotherm [60]. Similar bivariate skew-normal distributions were also identified in the Mg alloying systems in this study, as illustrated in Figure S35. A linear function was fitted to the distribution, with the slope having a similar magnitude ($10^{-4}$ 1/K) as that obtained for the Ni-based alloy systems [60]. Such significant scattering between $\Delta E_{\text{seg}}$ and $\Delta F_{\text{seg}}^{\text{vib}}$ leads to the unsatisfactory prediction of $\Delta F_{\text{seg}}^{\text{vib}}$ using the $\Delta E_{\text{seg}}$ feature. The prediction becomes even worse without including $\Delta E_{\text{seg}}$ in the feature sets, as $\Delta E_{\text{seg}}$ exhibits the highest correlation with $\Delta F_{\text{seg}}^{\text{vib}}$ among all features, as illustrated in the correlation matrices (Figures S4-9) and feature importance (Figure S15). The SOAP feature set, which is believed to comprehensively capture the structural features of the local atomic environment, fails to predict $\Delta F_{\text{seg}}^{\text{vib}}$, implying that structural features before solute segregation alone are insufficient for predicting vibrational entropy contribution of solute segregation. This could be interpreted as the change in bond stiffness, crucial in determining the variation of vibrational frequency when introducing a solute at a GB site, which cannot be precisely described using the structural features of the pre-segregation local atomic environment. Other energetic features describing the system before solute segregation, such as potential energy and $F^{\text{vib}}$, although still among the highest important features, exhibit much lower importance compared to $\Delta E_{\text{seg}}$ (Figure S15). Overall, these outcomes highlight that the features after segregation could be crucial in predicting the vibrational entropy contribution, with $\Delta E_{\text{seg}}$ being particularly important as it describes the energy difference before and after introducing the solute element. For better prediction of $\Delta F_{\text{seg}}^{\text{vib}}$, incorporating more structural and energetic features after solute segregation would be beneficial.
Despite the relatively low prediction accuracy of per-site $\Delta F_{\text{seg}}^{\text{vib}}$, the overall trend of its contribution to segregation behavior, specifically, the averaged solute concentration at GBs as calculated using the Langmuir-McLean model, is well captured. As illustrated in Figure S36, the solute concentration at 600 K, when computed using the predicted $\Delta F_{\text{seg}}$ values, deviates only slightly from the values obtained using the true $\Delta F_{\text{seg}}$ values (see Figure S36(b)). The ML model’s performance in predicting solute concentration at finite temperatures is comparable to its performance in predicting solute concentration at 0 K using $\Delta E_{\text{seg}}$ alone (see Figure S36(a)). This observation aligns with the relatively small absolute value of $\Delta F_{\text{seg}}^{\text{vib}}$ in comparison to segregation energy, reinforcing its limited impact on segregation trends. These results demonstrate the capability of the ML model in predicting solute concentration at finite temperatures, which serves as a crucial parameter for comparison with experimental measurements, as discussed in the next section.
@@
Figure 7: (a) Summary of experimentally measured concentrations of various solutes at GB versus bulk in Mg alloys [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. The concentrations of (b) Nd, (c) Ca, (d) Y, (e) Al, and (f) Zn solutes at GBs within the simulated Mg polycrystal as a function of their bulk solute concentrations, incorporating the vibrational entropy contribution $\Delta E_{\text{seg}}+\Delta F_{\text{seg}}^{\text{vib}}$ (represented by solid lines) and considering only the segregation energy $\Delta E_{\text{seg}}$ (represented by dashed lines) across various temperature regimes following the Langmuir–McLean model, as denoted in Equations (12) and (13). Experimental data presented in (a) are illustrated as black dots on the corresponding graphs, labeled with the first authors’ names and the thermomechanical processing temperatures of the Mg alloys.
@@
Experimental data of $X_{\text{GB}}$ versus $X_{\text{bulk}}$ are summarized in Figure 7(a). Notably, the Nd solute exhibits significantly higher $X_{\text{GB}}$ than other solute elements under similar $X_{\text{bulk}}$ and temperature conditions. This experimental observation is in good agreement with the simulation results presented in Figure 3, which demonstrate an elevated $X_{\text{GB}}$ for Nd at both 0.1 at.and 1 at.$X_{\text{bulk}}$ across the temperature range up to 800 K. This agreement between experimental and simulated data underscores Nd’s exceptional energetic preference for segregation at GBs, establishing it as the most susceptible among the studied alloying elements to undergo segregation. However, it is important to note that the simulation predicts the $X_{\text{GB}}$ of Nd solute with an overestimation by a factor of 4 to 6 in the temperature regime from 500 to 600 K (see Figure 7(b)). An overestimation in $X_{\text{GB}}$ for the Ca and Al solutes was also identified, with a factor of 2 to 3 across the relevant temperatures and $X_{\text{bulk}}$ regimes (see Figure 7(c,e)). A relatively good agreement was observed between experimental measurements and simulated predictions of $X_{\text{GB}}$ for Y and Zn solutes. The experimentally measured $X_{\text{GB}}$ values fall within the range predicted by atomistic simulations. To the best of our knowledge, no experimental segregation data on the chemical distribution of Li in GB were available. This is primarily attributed to the research focus on forming Mg-Li intermetallic phases to enhance the mechanical performances of Mg alloys, rather than on engineering GB properties below the solid solubility limit [79, 80].
@@
The overestimation of $X_{\text{GB}}$ could be attributed to factors related to both theoretical methodologies and experimental measurements, which are not regarded as inherent limitations of the approach. Firstly, the Langmuir–McLean model, employed in our current study, assumes an infinitesimally diluted solid solution, where solute-solute interactions are not accounted for. Experimental Mg alloy systems are often more complex than simple binary compositions, making it necessary to account for synergistic effects from solute-solute interactions. This simplification could potentially lead to disparities between experimental observations and theoretical predictions, e.g., as demonstrated in our recent APT study on the Mg-Zn-Ca alloy systems [36]. There, it was shown that the presence of Zn significantly reduced the concentration of Ca at GBs. This was evident from a decrease from 2.39 at.in a Mg-Ca binary system with $X_{\text{bulk}}$ of Ca at 0.163 at.to 1.95 at.in a Mg-Zn-Ca ternary system with even higher bulk Ca concentration of 0.421 at.(see Mouhib, 673 K in Figure 7(c)) [36]. Given that this binary alloy system for Ca segregation is the only example available in the literature, it is noteworthy that the experimentally measured $X_{\text{GB}}$ for this system demonstrates a closer alignment with theoretical predictions compared to results from ternary or quaternary alloy systems. Moreover, the available experimental data on Y segregation at GBs were all obtained from Mg-Y binary alloy systems (Figure 7(d)), demonstrating a good correlation with theoretical predictions. To investigate solute segregation behavior at higher concentrations, hybrid Monte-Carlo/molecular dynamics (MC/MD) approaches can be employed to provide thermodynamic insights into solute clustering, precipitation, and the formation of intermetallic phases at GBs [81, 82, 83].
@@
Additionally, the discrepancy may stem from the limitations in experimental measurements, which were performed on selected individual GBs, contrasting with atomistic simulations that encompassed a polycrystal with over 38,000 GB sites in a GB network with various macro and microscopic characteristics. As revealed in our recent APT study, significant variations were observed in the segregation behavior of Nd among GBs in Mg alloys, which possess distinct macroscopic characteristics [35]. Notably, the concentration of Nd at GBs ($X_{\text{GB}}$) was found to be more than twice as high in a GB with a misorientation angle of 61.1 compared to one with an angle of 35.8. This scattering in GB segregation can primarily be attributed to variations in the local atomic environments, indicating the crucial role of microscopic characteristics in influencing solute distribution at the atomic level. Furthermore, in experimental conditions, there is always competition with other defects, such as dislocations or stacking faults, which cannot be switched off. If a large number of solutes get trapped in these defects, this can also lead to less solute being available to segregate at the GBs. This underscores the need for defect phase diagrams [84] to better understand solute distribution mechanisms under different chemical potential conditions.
@@
Furthermore, artifacts and approximations in both experimental and simulation methodologies can significantly impact the outcome solute segregation behavior. For instance, during the post-processing of APT and EDS data, assumptions about the GB thickness can lead to notable discrepancies in the measurements of $X_{\text{GB}}$. Similarly, the definition of GB sites according to structural analysis algorithms for segregation energy calculations influences the segregation energy spectrum and, consequently, the estimated solute concentration at GBs. In atomistic simulations, the choice of methodological assumptions for calculating $\Delta F_{\text{seg}}^{\text{vib}}$, such as those made by the HA method, can influence results. Specifically, the HA method’s inability to capture anharmonic behaviors—–like structural transitions—–and its omission of other significant entropy contributions, especially at elevated temperatures, are notable limitations of this model. The choice of interatomic potential also impacts the simulation outcomes. For instance, notable differences in segregation energy spectra were observed when comparing results obtained using the MEAM [51] and EAM [85] potentials, see Figure S37. Nevertheless, both potentials yielded similar mean segregation energy values of approximately -0.08 eV/atom for Al, indicating a comparable overall tendency of Al solutes to segregate to GBs. The limitations of semi-empirical potentials in accurately capturing the variation in solute segregation behaviors across the diverse local atomic environments of GBs and a broad range of alloy systems have been reported [86]. Hybrid multiscale modeling approaches, such as quantum mechanical/molecular mechanical (QM/MM), could offer QM-level accurate results while circumventing the size constraints typically associated with ab-initio calculations.
IV.3 Generalizability of ML models
[…]
Moreover, previous ML models for solute segregation in Mg GBs were often limited to single elements, such as Al [41] or Y [42], or used different interatomic potentials for the Mg matrix when handling multiple elements [40]. In contrast, this work considers the segregation behavior of multiple solutes simultaneously within a unified framework, which allows us to consider the effect of intrinsic properties of the solute elements on segregation tendencies. Segregation behavior at GBs is not only influenced by the characteristics of GB sites but also by the intrinsic properties of solutes, including atomic size, electronegativity, and bonding characteristics. Our findings highlight the variability in segregation energy spectra and feature importance rankings among the studied solutes, reflecting the complex interplay between solutes’ intrinsic properties and the local atomic environment at GB sites.
The ML models developed in this work utilize a combination of energetic and structural descriptors of the Mg matrix to predict segregation behavior. While these models were trained and validated using these six solute elements, their applicability to other elements in Mg alloys depends on the generalizability of the features used. The energetic and structural features, such as hydrostatic stress, atomic volume, and potential energy, serve as general descriptors that capture fundamental aspects of segregation behavior influenced by the characteristics of GB sites and are transferable across different solutes. Incorporating additional intrinsic features of solutes, such as misfit volume, cohesive energy, electronegativity differences, and diffusion coefficients, could further enhance the model’s ability to predict segregation behavior for a broader range of solutes by providing deeper insights into how solutes interact with the Mg matrix and local GB environments.
Furthermore, expanding the training dataset to include a wider range of solute elements and more diverse local atomic environments is critical for improving the models’ robustness and generalizability. Performing cross-validation with experimental or computational data for additional solutes will ensure that the models capture broader trends and remain reliable for new predictions. These considerations underscore the potential for extending the current ML framework to predict segregation behavior across a broader spectrum of solute elements, ultimately providing a more comprehensive tool for understanding and engineering GB segregation in Mg alloys.
@@ V Conclusions
In this study, the GB segregation behavior of Nd, Ca, Y, Li, Al, and Zn solute elements in polycrystalline Mg at 0 K and finite temperatures was systematically investigated using atomistic simulations. Machine learning (ML) models were employed to predict the segregation energy and free energy, incorporating energetic and structural descriptors. The GB solute concentrations in these Mg alloy systems were calculated following the Langmuir–McLean isotherm and compared with experimentally measured local chemical distributions at various temperature and bulk concentration regimes. Key insights from the study are summarized as follows:
@@
Author contributions
Z.X. performed atomistic simulations, applied machine learning models, and wrote the original draft. A.A. performed atomistic simulations and constructed local atomic environment descriptors. U.K. established the machine learning workflow. All authors contributed to the discussion and editing of the final manuscript.